# ResQuNNs:Towards Enabling Deep Learning in Quantum Convolution Neural Networks

### v5

@@ Abstract
This paper focuses on methidenhances Quantum Convolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and addressing the challenges in training deep QuNNs. Traditional QuNNs rely on static quanvolutional layers, limiting their feature extraction capabilities. Our approach enables the training of these layers, significantly improving the scalability and learning potential of QuNNs. However, deep QuNNs face difficulties in gradient-based optimization due to limited gradient flow across multiple layers. To overcome this, we propose Residual Quanvolutional Neural Networks (ResQuNNs), which utilize residual learning by adding skip connections between quanvolutional layers. These residual blocks enhance gradient flow throughout the network, facilitating effective training in deep QuNNs. Moreover, We provide empirical evidence on the optimal placement of these residual blocks, demonstrating how strategic configurations improve gradient flow and lead to more efficient training. Our findings represent a significant advancement in quantum deep learning, opening new possibilities for both theoretical exploration and practical quantum computing applications.
@@ 1 Introduction
Quantum Convolutional Neural Networks (QCNNs) represent an innovative fusion of quantum computing principles with traditional convolutional neural network (CNN) architectures [1, 2]. These hybrid classical-quantum models utilize parameterized quantum circuits (PQCs) a.k.a ansatz, composed of parameterized rotation gates such as $R_{x}$, $R_{y}$, or $R_{z}$, functioning analogously to convolutional and pooling operations in classical CNNs [3, 4]. The optimization, however, is performed on classical computers [5, 6]. A notable advantage of QCNNs lies in their ability to harness the vast Hilbert space offered by quantum mechanics, which surpasses the capabilities of classical CNNs [7]. This feature allows QCNNs to capture spatial relationships in image data more effectively. Furthermore, the potential future inclusion of higher-level quantum systems, such as qutrits or ququads [8, 9], promises even more sophisticated image comprehension capabilities.
There are two primary variants of QCNNs, each leveraging quantum mechanics principles like superposition, entanglement, and interference. The first type, Quantum-Inspired CNNs, proposed in [3], mirrors the structure of classical CNNs but replaces convolutional and pooling layers with deeper PQCs. This design demands higher qubit coherence to fully exploit quantum parallelism and entanglement, posing a challenge for current quantum hardware. Nevertheless, a significant research has been dedicated to explore this very architectural framework for various applications [5, 10]. The second type, Quanvolutional Neural Networks (QuNNs), first introduced in [11], and visually depicted in Figure 1(a), focuses on quantum-enhanced convolutional layers. These networks encode classical image pixels into quantum states, apply unitary transformations via PQCs, and then measure the qubits to create output image channels, which are subsequently processed by classical fully connected layers. This approach offers a more feasible implementation with current quantum technology. Therefore, in this paper, we also delve into QuNNs with a focus on identifying the challenges faced in training QuNNs and proposing effective solutions to address these issues.
@@
The QuNN architecture depicted in Figure 1, widely adopted in state-of-the-art research, is extensively utilized, for various applications such as 2D and 3D radiological image classification [12], high energy physics data analysis [13], and object detection and classification [14, 15]. In these contexts, the quantum circuit in QuNNs primarily functions as a feature extractor and is not trained during the quanvolution operation. Consequently, the learning process is primarily driven by the subsequent classical network. Although there are recent studies that employ trainable quantum circuits in quanvolutional layers, to the best of our knowledge, no existing literature addresses multi-layered QuNNs. Table 1 some summarizes recent works utilizing QuNNs with both trainable and non-trainable quanvolutional layers. To achieve full scalability in QuNNs, it is crucial to ensure the trainability of underlying quanvolutional layers in multi-layered architectures.
@@ 1.1 Motivational Case Study
Figure 3: Gradients Propagation in QuNNs for Two (upper pannel) and Three (lower panel) Quanvolutional Layers.
While a single trainable quanvolutional layer enhances QuNNs’ training efficacy, as depicted in Figure 2, scalability of QuNNs requires multiple trainable layers. However, challenges emerge when multiple quanvolutional layers are trained concurrently, as gradients are exclusively accessible only in the last quanvolutional layer, thus limiting the optimization process only to this layer. Figure 3 demonstrates this by showing gradients for two and three trainable quanvolutional layers. We observe that the gradients of only the final quanvolutional layer are accessible, while those of the preceding layers are effectively None. The restricted gradient flow means that in multi-layered QuNNs, only the last quanvolutional layer contributes in the learning process, leaving earlier layers remain uninvolved in optimization. This presents a significant challenge in effectively training multi-layered QuNNs, thereby necessitating innovative approaches to address this challenge for scalable and effective QuNN architectures.
@@ 1.2 Our Contributions
  • Both Classical and Quantum Postprocessing. State-of-the-art QuNNs utilize classical networks to postprocess the output of quanvolutional layers. However, to demonstrate the effectiveness of the residual approach in multi-layered QuNNs (with no influence on the overall learning from classical layers), other than classical postprocessing layer, we employ the quantum circuit itself for postprocessing the results of the quanvolutional layers.
@@ 3 ResQuNN Methodology
    Encoding Classical Data into Quantum States. We use a kernel of size $2\times 2$ which slides over the entire input image with stride 2. This results in $4$ classical features extracted after every step, i.e., $x=\{x_{1},x_{2},x_{3},x_{4}\}$. These features are then encoded into quantum states, specifically into the rotation angles of $RY$ gates. The initial state of each qubit |$\ket{q_{i}}$ is the ground state |$\ket{0}$, and the RY gate applies a rotation by $x_{i}$ radians on the Bloch sphere, which can be described by the following Equation:
    $U_{{\text{Enc}}_{1}}=|q_{i}\rangle=RY(x_{i})|0\rangle=\cos\left(\frac{x_{i}}{2}\right)|0\rangle+\sin\left(\frac{x_{i}}{2}\right)|1\rangle$ (1)
@@
    $U_{\text{Rotation}}=RX(\theta_{i})|q_{i}\rangle=\cos\left(\frac{\theta_{i}}{2}\right)|q_{i}\rangle-i\sin\left(\frac{\theta_{i}}{2}\right)|q_{i}\rangle$ (3)
@@
    To introduce entanglement between qubits in the quanvolutional layers, CNOT gates are applied between nearest neighbor qubits. The entanglement operation in this context can be defined as:
@@
    In case of two quanvolutional layers, the first quanvolutional layer (QCL1) has a depth of $4$, while that of second (QCL2) has a depth of $1$. This configuration is strategically chosen to distinctly evaluate the impact on performance when gradients from both layers are accessible in ResQuNNs, in contrast to situations where the gradients of only the second layer are accessible. It is noteworthy that at reduced and/or equal depths in both layers, the performance difference may not be as pronounced or readily observable even if the gradients in second layer are also available. Applied to the initial state $|0000\rangle$, the final state after the first QCL can be expressed as:
@@
    $U_{{\text{Enc}}_{2}}=|q_{i}\rangle=RY(y_{i})|0\rangle=\cos\left(\frac{y_{i}}{2}\right)|0\rangle+\sin\left(\frac{y_{i}}{2}\right)|1\rangle$ (10)
@@
    |$\ket{\psi_{\text{QCL}_{2}}}=U_{\text{circuit}_{2}}\ket{y_{i}}$ (13)
@@ 3.1.1 Classical layer to postprocess quanvolutional layer’s output
    Let’s assume the output of the quanvolutional layer is a set of measurements $\{M=m_{1},m_{2}\ldots m_{n}\}$, where each $m_{i}$ represents the measurement of a qubit in terms of classical information, such as the expectation values in computational basis. This result is transformed to a vector format and then passed to the fully connected layer with $10$ neurons. The output feature size from quanvolutional layers is different and depends on the kernel size and stride similar to that of classical convolutional layers. We experiment with a kernel size of $2\times 2$ and stride of $2$. The output feature size after each quanvolutional layer in case of different residual configuration (label 5 in Figure 5) is presented in the Table 2.
@@
    • $f$ is the activation function (SoftMax in our case).
@@ 3.1.2 Quantum Circuit to postprocess quanvolutional layer’s output
    We also use quantum circuit to postprocess the result from quanvolutional layer, as highlighted by label 3 in Figure 5. This is primarily for two reasons: firstly to demonstrate the working of fully quantum QuNNs, and secondly, to better understand the importance of residual connections by removing the influence of classical layers being used at the end.
    The use of a quantum circuit for post-processing introduces certain computational overheads, mainly due to the prerequisite of re-encoding the (classical) outputs from the quanvolutional layers, a process that entails considerable computational expense. Also the following trainable quantum circuit for postprocessing quanvolutional layer’s results after the encoding takes longer to produce the results than a simple classical neuron layer. We use $4$ classes from MNIST dataset for experimental simplification. For the purpose of encoding the outputs from the quanvolutional layer, Amplitude Encoding [26], is used which can be described by the following Equation:
@@
    where $x_{i}$ is the $i^{th}$ element of $x$ and |$\ket{i}$ is the $i^{th}$ computational basis state. We used amplitude encoding for postprocessing the results of quanvolutional layers primarily because it allows to encode $2^{n}$ features in $n$ qubits. Based on Table 2, the smallest output feature size from quanvolutional layer(s) is $7\times 7\times 1=49$, which requires atleast $49$ qubits via angle encoding (one qubit per feature) and only $6$ qubits via amplitude encoding. Similarly, the maximum output size of quanvolutional layer is $28\times 28\times 1=784$, requiring $784$ qubits, if encoded in qubit rotation angles, whereas Amplitude Encoding reduces the requirement to merely $10$ qubits. The use of $784$ qubits is not practical in current NISQ devices, therefore we use amplitude encoding to encode the qunavolutional layer’s result for further processing by a quantum circuit. To ensure a consistent impact of the quantum circuit’s post-processing across all residual configurations, a $10$-qubit circuit is utilized for processing the outputs of the quanvolutional layer, corresponding to the maximum feature size of $784$. However, as mentioned earlier, we use $4$ classes, therefore, only $4$ qubits are measured each corresponding to one class. The output of quanvolutional, once encoded, is subjected to the unitary evolution through a simple quantum circuit, which can be described by the following Equation:
@@ 5.1.1 Analysis of Gradients Accessibility
In the case of No Residual setting, as previously shown in Figure 3, we have already demonstrated the fact that the gradients are exclusively accessible for the second quanvolutional layer. Therefore, we now proceed to investigate the gradient propagation through the quanvolutional layers for all other residual settings (from label 5 in Figure 5), the results of which are illustrated in Figure 7. It is worth mentioning here that while the gradient accessibility results are presented with quantum layer postprocessing, the accessibility of gradients remains the same across layers with classical postprocessing as well.
Figure 7: Gradients Propagation Through the Network for Two Quanvolutional Layers for Different Residual Configurations. X=Input, O1, O2=Output of first and second quanvolutional layer respectively. To demonstrate gradients accessibility, we consider each quanvolutional layer with a quantum circuit depth of $1$ since gradient propagation is not effectred by depth.
@@ 5.1.2 Analysis of Training Performance
    Figure 8: Two Quanvolutional Layers - Training results of all residual configurations with classical postprocessing and $\textbf{2}\times\textbf{2}$ Kernel.
@@
    Figure 9: Two Quanvolutional Layers - Comparison of ResQuNNs with benchmark models for $\textbf{2}\times\textbf{2}$ Kernel.
@@
    Figure 10: Two Quanvolutional Layers - Training results of all residual configurations with quantum postprocessing $\textbf{2}\times\textbf{2}$ Kernel.
@@ 5.2 Three Quanvolutional Layers
Figure 11: Gradients Propagation Through the Network for Three Quanvolutional Layers for (O1+O2)+O3 (top) and ((X+O1)+O2)+O3 (bottom)
@@ 6 Conclusion
In this paper, we presented a novel framework for developing Quanvolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and the concept of Residual Quanvolutional Neural Networks (ResQuNNs) for enabling the deep learning in QuNNs. Our approach addresses the significant challenges of gradient flow and adaptability in deep QuNN architectures, enabling enhanced training performance and scalability. A key contribution of this work is the development of trainable quanvolutional layers, which enhance the learning potential of QuNNs, and the strategic incorporation of residual blocks to ensure comprehensive gradient accessibility across the network in multi-layered QuNNs, thus, enabling deeper, more complex QuNNs architectures to be effectively trained. Our empirical findings highlight the importance of optimal placement of residual blocks, providing critical insights for future QuNN architecture designs with even more deeper layers. In summary, our work signifies a substantial leap in quantum deep learning, offering new directions for theoretical and practical quantum computing research. By tackling core issues of adaptability and gradient flow, our work paves the way for the development of more robust and efficient quantum neural networks, marking a pivotal step toward harnessing the full potential of quantum technologies.

### v6

# Deep Quanvolutional Neural Networks with Enhanced Trainability and Gradient Propagation

@@ Abstract
In this paper, we explore methods to enhance the performance of one of the frequently used variants of Quantum Convolutional Neural Networks, known as Quanvolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and addressing the challenges associated with training multi-layered or deep QuNNs. Traditional QuNNs mostly rely on static (non-trainable) quanvolutional layers, limiting their feature extraction capabilities. Our approach enables the training of these layers, significantly improving the scalability and learning potential of QuNNs. However, multi-layered deep QuNNs face difficulties in gradient-based optimization due to limited gradient flow across all the layers of the network. To overcome this, we propose Residual Quanvolutional Neural Networks (ResQuNNs), which utilize residual learning by adding skip connections between quanvolutional layers. These residual blocks enhance gradient flow throughout the network, facilitating effective training in deep QuNNs, thus enabling deep learning in QuNNs. Moreover, we provide empirical evidence on the optimal placement of these residual blocks, demonstrating how strategic configurations improve gradient flow and lead to more efficient training. Our findings represent a significant advancement in quantum deep learning, opening new possibilities for both theoretical exploration and practical quantum computing applications.
@@ 1 Introduction
Quantum Machine Learning (QML) combines quantum computing with machine learning to process and analyze data using quantum algorithms and circuits[1]. By leveraging quantum phenomena such as superposition and entanglement, QML aims to provide computational advantages over classical methods, especially for complex, high-dimensional problems[2, 3, 4]. Quantum Convolutional Neural Networks (QCNNs) represent an innovative fusion of quantum computing principles with traditional convolutional neural network (CNN) architectures [5, 6]. These hybrid classical-quantum models utilize parameterized quantum circuits (PQCs) a.k.a ansatz, composed of parameterized rotation gates such as $R_{x}$, $R_{y}$, or $R_{z}$, functioning analogously to convolutional and pooling operations in classical CNNs [7, 8]. The optimization, however, is performed on classical computers [9, 10, 11, 12]. A notable advantage of QCNNs lies in their ability to harness the vast Hilbert space offered by quantum mechanics, which surpasses the capabilities of classical CNNs [13]. This feature allows QCNNs to capture spatial relationships in image data more effectively. Furthermore, the potential future inclusion of higher-level quantum systems, such as qutrits or ququads [14, 15], promises even more sophisticated image comprehension capabilities.
There are two primary variants of QCNNs, each leveraging quantum mechanics principles like superposition, entanglement, and interference. The first type, Quantum-Inspired CNNs, proposed in [7], mirrors the structure of classical CNNs but replaces convolutional and pooling layers with deeper PQCs. This design demands higher qubit coherence to fully exploit quantum parallelism and entanglement, posing a challenge for current quantum hardware. Nevertheless, a significant research has been dedicated to explore this very architectural framework for various applications [9, 16]. The second type, Quanvolutional Neural Networks (QuNNs), first introduced in [17], and visually depicted in Figure 1(a), focuses on quantum-enhanced convolutional layers. These networks encode classical image pixels into quantum states, apply unitary transformations via PQCs, and then measure the qubits to create output image channels, which are subsequently processed by classical fully connected layers. This approach offers a more feasible implementation with current quantum technology. Hence, we also delve into QuNNs with a focus on identifying the challenges faced in training QuNNs and proposing effective solutions to address these issues.
@@
The QuNN architecture depicted in Figure 1, widely adopted in state-of-the-art research for various applications such as 2D and 3D radiological image classification [18], high energy physics data analysis [19], and object detection and classification [20, 21]. In these contexts, the quantum circuit in QuNNs primarily functions as a feature extractor and is not trained during the quanvolution operation. Consequently, the learning process is primarily driven by the subsequent classical network. Although there are recent studies that employ trainable quantum circuits in quanvolutional layers, to the best of our knowledge, no existing literature addresses multi-layered QuNNs. Table 1 some summarizes recent works utilizing QuNNs with both trainable and non-trainable quanvolutional layers. To achieve full scalability in QuNNs, it is crucial to ensure the trainability of underlying quanvolutional layers in multi-layered architectures.
@@ 1.1 Motivational Case Study
Figure 3: Gradients Propagation in multi-layered QuNNs with Two (upper pannel) and Three (lower panel) Quanvolutional Layers.
While a single trainable quanvolutional layer enhances QuNNs’ training efficacy, as depicted in Figure 2, scalability of QuNNs requires multiple trainable layers. However, challenges emerge when multiple quanvolutional layers are trained concurrently, as gradients are exclusively accessible only in the last quanvolutional layer, thus limiting the optimization process only to this layer. The gradient accessibility in multi-layered QuNNs is hindered because each quantum layer performs qubit measurement, thus collapsing the quantum state and breaking the chain of differentiability. Figure 3 demonstrates this by showing gradients for two and three trainable quanvolutional layers. We observe that the gradients of only the final quanvolutional layer are accessible, while those of the preceding layers are effectively None. The restricted gradient flow means that in multi-layered QuNNs, only the last quanvolutional layer contributes in the learning process, leaving earlier layers remain uninvolved in optimization. This presents a significant challenge in effectively training multi-layered QuNNs, thereby necessitating innovative approaches to address this challenge for scalable and effective QuNN architectures.
@@ 1.2 Our Contributions
  • Classical and Quantum Postprocessing. State-of-the-art QuNNs utilize classical networks to postprocess the output of quanvolutional layers. However, to demonstrate the effectiveness of the residual approach in multi-layered QuNNs (with no influence on the overall learning from classical layers), other than classical postprocessing layer, we employ the quantum circuit itself for postprocessing the results of the quanvolutional layers.
@@ 3 ResQuNN Methodology
    Encoding Classical Data into Quantum States. We use a kernel of size $2\times 2$ which slides over the entire input image with stride 2. This results in $4$ classical features extracted after every step, i.e., $x=\{x_{1},x_{2},x_{3},x_{4}\}$. These features are then encoded into quantum states, specifically into the rotation angles of $RY$ gates. This approach is commonly called as angle encoding [32]. Each qubit |$\ket{q_{i}}$ is prepared in the ground state |$\ket{0}$, and the $RY$ gate applies a rotation by $x_{i}$ radians on the Bloch sphere, which can be described by the following Equation:
    $U_{{\text{Enc}}_{1}}=|q_{i}\rangle=RY(x_{i})|0\rangle\quad\text{where}\quad RY(\theta)=\begin{pmatrix}\cos(x/2)&-\sin(x/2)\\ \sin(x/2)&\cos(x/2)\end{pmatrix}$ (1)
@@
    $U_{\text{Rotation}}=RX(\theta_{i})|q_{i}\rangle\quad\text{where}\quad RX(\theta)=\begin{pmatrix}\cos(\theta/2)&-i\sin(\theta/2)\\ -i\sin(\theta/2)&\cos(\theta/2)\end{pmatrix},$ (3)
@@
    To introduce entanglement between qubits in the quanvolutional layers, CNOT gates are applied between nearest neighbor qubits. Since we use total of $4$ qubits, the entanglement operation in this context can be defined as:
@@
    In the proposed architecture, the first quanvolutional layer (QCL1) is designed with a circuit depth of $4$, while the second quanvolutional layer (QCL2) exhibits a reduced depth of $1$. Both QCL1 and QCL2 incorporate identical quantum operations, consisting of single-qubit parameterized rotation gates and two-qubit CNOT entangling gates, as illustrated in label 1 of Figure 5. The use of only single and two-qubits gates makes the quantum circuits compatible with NISQ devices. The distinction between the two layers, i.e., QCL1 and QCL2 lies solely in the number of repetitions of the quantum circuit, thereby differentiating their respective depths. This configuration (deep quantum layers in QCL1 and shallow layers in QCL2) is strategically chosen to distinctly evaluate the impact on performance when gradients from both layers are accessible in ResQuNNs, compared to the scenarios where the gradients of only the second layer are accessible. The reason is that when QCL1 is configured with a reduced depth, or when both QCL1 and QCL2 have equal depths, the effect of QCL1 on resulting performance difference may not be as pronounced or readily observable, even in cases where gradients from the second layer are accessible. Applied to the initial state $|0000\rangle$, the final state after the first QCL can be expressed as:
@@
    $U_{{\text{Enc}}_{2}}=|q_{i}\rangle=RY(y_{i})|0\rangle$ (10)
@@
    |$\ket{\psi_{\text{QCL}_{2}}}=U_{\text{circuit}_{2}}\ket{y_{\text{enc}}}$ (13)
@@ 3.1.1 Classical layer to postprocess quanvolutional layer’s output
    Let’s assume the output of the quanvolutional layer is a set of measurements $\{M=m_{1},m_{2}\ldots m_{n}\}$, where each $m_{i}$ represents the measurement of a qubit in terms of classical information, such as the expectation values in computational basis. This result is transformed to a vector format and then passed to the fully connected layer with $10$ neurons. The output feature size from quanvolutional layers is different and depends on the kernel size and stride similar to that of classical convolutional layers. We experiment with a kernel size of $2\times 2$ and stride of $2$. The output feature size after each quanvolutional layer for different residual configuration (label 5 in Figure 5) is presented in the Table 2.
@@
    • $f$ is the SoftMax activation function, which converts a vector of K real numbers into a probability distribution of K possible outcomes. It is frequently used in multiclass classification problems.
@@ 3.1.2 Quantum Circuit to postprocess quanvolutional layer’s output
    We also use quantum circuit to postprocess the result from quanvolutional layer, as highlighted by label 3 in Figure 5. This is primarily for two reasons: firstly to demonstrate the working of fully quantum QuNNs, and secondly, to better understand the importance of residual connections by removing the influence of classical layers being used at the end. The use of a quantum circuit for post-processing introduces certain computational overheads, mainly due to the prerequisite of re-encoding the (classical) outputs from the quanvolutional layers, a process that entails considerable computational expense. Also, the following trainable quantum circuit for postprocessing quanvolutional layer’s results after the encoding, takes longer to produce the results than a simple classical neuron layer. We use $4$ classes from MNIST dataset for experimental simplification. For the purpose of encoding the outputs from the quanvolutional layer, Amplitude Encoding [33], is used which can be described by the following Equation:
@@
    where $x_{i}$ is the $i^{th}$ element of $x$ and |$\ket{i}$ is the $i^{th}$ computational basis state. We used amplitude encoding for postprocessing the results of quanvolutional layers primarily because it allows to encode $2^{n}$ features in $n$ qubits [32]. Based on Table 2, the smallest output feature size from quanvolutional layer(s) is $7\times 7\times 1=49$, which requires atleast $49$ qubits via angle encoding (one qubit per feature) and only $6$ qubits via amplitude encoding. Similarly, across the range of experiments we performed, the maximum output size of quanvolutional layer is $28\times 28\times 1=784$, requiring $784$ qubits, if encoded in qubit rotation angles, whereas amplitude encoding reduces the requirement to merely $10$ qubits. The use of $784$ qubits is not practical in current NISQ devices, therefore we use amplitude encoding to encode the qunavolutional layer’s result for further processing by a quantum circuit. To ensure a consistent impact of the quantum circuit’s post-processing across all residual configurations, a $10$-qubit circuit is utilized for processing the outputs of the quanvolutional layer, corresponding to the maximum feature size of $784$. However, as mentioned earlier, we use $4$ classes, therefore, only $4$ qubits are measured each corresponding to one class. The output of quanvolutional, once encoded, is subjected to the unitary evolution through a simple quantum circuit, which can be described by the following Equation:
@@ 5.1.1 Analysis of Gradients Accessibility
In the case of No Residual setting (which refers to the configuration in which no residual blocks are incorporated within the multi-layered QuNNs and serves as a benchmark for comparison) we have already demonstrated in Figure 3 that the gradients are exclusively accessible for the second quanvolutional layer. Therefore, we now proceed to investigate the gradient propagation through the quanvolutional layers for all other residual settings (from label 5 in Figure 5), the results of which are illustrated in Figure 7. It is worth mentioning here that while the gradient accessibility results are presented with quantum layer postprocessing, the accessibility of gradients remains the same across layers with classical postprocessing as well.
Figure 7: Gradients Propagation Through the Network for Two Quanvolutional Layers for Different Residual Configurations. X=Input, O1, O2=Output of first and second quanvolutional layer respectively.
@@ 5.1.2 Analysis of Training Performance
    Figure 8: Two Quanvolutional Layers - Training results of all residual configurations with classical postprocessing and $2\times 2$ Kernel. No Residual setting refers to multi-layered QuNNs without any residual block
@@
    Figure 9: Two Quanvolutional Layers - Comparison of ResQuNNs with benchmark models for $2\times 2$ Kernel. No Residual setting refers to multi-layered QuNNs without any residual block
@@
    Figure 10: Two Quanvolutional Layers - Training results of all residual configurations with quantum postprocessing ×$\times 2$ Kernel. No Residual setting refers to multi-layered QuNNs without any residual block
@@ 5.2 Three Quanvolutional Layers
Figure 11: Gradients Propagation Through the Network for Three Quanvolutional Layers for (O1+O2)+O3 (left) and ((X+O1)+O2)+O3 (right)
@@ 6 Conclusion
In this paper, we presented a framework for developing Quanvolutional Neural Networks (QuNNs) by introducing trainable quanvolutional layers and the concept of Residual Quanvolutional Neural Networks (ResQuNNs) for enabling the deep learning in QuNNs. Our approach addresses the significant challenges of gradient flow and adaptability in deep QuNN architectures, enabling enhanced training performance and scalability. A key contribution of this work is the development of trainable quanvolutional layers, which enhance the learning potential of QuNNs, and the strategic incorporation of residual blocks to ensure comprehensive gradient accessibility across the network in multi-layered QuNNs, thus, enabling deeper, more complex QuNNs architectures to be effectively trained. Our empirical findings highlight the importance of optimal placement of residual blocks, providing critical insights for future QuNN architecture designs with even more deeper layers. In summary, our work signifies a substantial leap in quantum deep learning, offering new directions for theoretical and practical quantum computing research. By tackling core issues of adaptability and gradient flow, our work paves the way for the development of more robust and efficient quantum neural networks, marking a pivotal step toward harnessing the full potential of quantum technologies.
Data availability
The datasets used and/or analysed during the current study available from the corresponding author on reasonable request.